OpenStudy (anonymous):

Help for a medal and fan?!?! Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

Hi adore

OpenStudy (anonymous):

Solve for y in the first equation so you have\[y=2x+1\] Then, you substitute that in for y in the second equation, then you have \[-4x+(2x+1)=-1\] Solve it, then once you find x, put that value in the first equation and solve.

OpenStudy (anonymous):

So how would I do this? Would I multiply the -4x and the 2x?

OpenStudy (anonymous):

Nope you would add.

OpenStudy (anonymous):

Oh, okay. What about the 1 and -1?

OpenStudy (anonymous):

Once you add. you will have -2x+1=-1. Then solve a two-step equation.

OpenStudy (anonymous):

-2x=-2?

OpenStudy (anonymous):

Yeppp. And x=1. So plus t

OpenStudy (anonymous):

Plug that 1 into X in the first equation to get your Y. c:

OpenStudy (anonymous):

Okay, I can finish this problem from here. Can you help with 1 more?

OpenStudy (anonymous):

Sure xD

OpenStudy (anonymous):

Thank you! Solve the system of equations. x + 3y = -1 2x + 2y = 6

OpenStudy (anonymous):

So the same thing ;o Solve for x in the first equation and get \[x=-3y-1\] Plug that into x in the second equation. \[2(-3y-1)+2y=6\] Solve for y, then plug that into an equation.

OpenStudy (anonymous):

So you multiply the 2 by everything in the parenthesis?

OpenStudy (anonymous):

Yess.

OpenStudy (anonymous):

-6y-2+2y=6.

OpenStudy (anonymous):

So, (-6y-2)+2y=6. So then you subtract the -6y by 2y to get -8y ?

OpenStudy (anonymous):

you add. and get -4y c:

OpenStudy (anonymous):

Close enough :/

OpenStudy (anonymous):

Then you add the 2 to the 6 to get: -4y=8?

OpenStudy (anonymous):

Yes! Yaaaaaay.

OpenStudy (anonymous):

Then you divide the 8 by -4y to get -2?

OpenStudy (anonymous):

Yess.

OpenStudy (anonymous):

Okay, I got it from here. Thank you so much for your help and time!

OpenStudy (anonymous):

Your welcome!